OpenStudy (anonymous):

A triangular prism has a volume of 5 m3. If each dimension of the prism is tripled, what is the new volume of the prism? Answer 135 m3 45 m3 15 m3 405 m3

OpenStudy (anonymous):

i got 45 m3

OpenStudy (ash2326):

Volume of a triangular prism= Area of the base \(\times\) height \[V=\sqrt {3} \frac{a^2}{2} \times h\] a= edge of the base triangle h= height of prism \[5=\sqrt {3} \frac{a^2}{2} \times h\] Now each dimension is tripled new edge= 3a new height=3h so \[V=\sqrt {3} \frac{(3a)^2}{2} \times 3h\] we get \[V=9 \times 3 \sqrt {3} \frac{(a)^2}{2} \times h\] \[V=27 \times \sqrt {3} \frac{(a)^2}{2} \times h\] We know \[5=\sqrt {3} \frac{a^2}{2} \times h\] so \[V=27*5=135 m^3\]

OpenStudy (anonymous):

tysm!